Forrester Company is considering buying new equipment that would increase monthly fixed costs from $123,000 to $187,000 and would decrease the current variable costs of $80 by $25 per unit. The selling price of $110 is not expected to change. Forrester's current break-even sales are $451,000 and current break-even units are 4,100. If Forrester purchases this new equipment, the revised break-even point in units would:

Increase by 700.

Increase by 820.

Increase by 4,920.

Decrease by 820.

Decrease by 700.
Answer #1

New variable cost=80-25=$55 per unit

Contribution margin=Sales-Variable cost  

=(110-55)=$55 per unit

Breakeven=Fixed expenses/Contribution margin

=(187000/55)=3400 units

Hence decrease in break even units

=4100-3400=700 units